OpenStudy (anonymous):

Solving for 3 variables questions :o help! Solve the system by elimination 1: x+5y-4z=-10, 2x-y+5z=-9, 2x-10y-5z=0 . 2. Solve the system by substitution: 2x-y+z=-4, z=5, -2x+3y-z=-10. 3. A food store makes a 11–pound mixture of peanuts, almonds, and raisins. The cost of peanuts is $1.50 per pound, almonds cost $3.00 per pound, and raisins cost $1.50 per pound. The mixture calls for twice as many peanuts as almonds. The total cost of the mixture is $21.00. How much of each ingredient did the store use?

OpenStudy (anonymous):

options for the 1st one are : (5, –1, 0) (–5, 1, 0) (–5, –1, 0) (–5, –1, –2) Options for the second are: (–8, 7, 5) (–8, –7, 5) (8, –7, 5) (–8, –7, –5) And the 3rd : 3 lbs peanuts, 6 lbs almonds, 2 lbs raisins 8 lbs peanuts, 1 lb almonds, 2 lbs raisins 6 lbs peanuts, 3 lbs almonds, 2 lbs raisins 8 lbs peanuts, 2 lbs almonds, 1 lbs raisins

OpenStudy (anonymous):

I guess the 2nd one technically gives you z so haha that's nice eh?

OpenStudy (anonymous):

ok I can try to do this but it's gonna take a while. do you want the steps or just the answers?

OpenStudy (anonymous):

Well If it's not too much to ask of you I wouldn't mind having the steps as well :o

OpenStudy (anonymous):

ok I'll give it a shot. give me a minute

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

still working on it :)

OpenStudy (anonymous):

Haha okay ^-^ Thanks for your time!

OpenStudy (anonymous):

I have the first two questions answered but the 3rd one is confusing me.

OpenStudy (anonymous):

Aw well any guess of yours on that one is as good as a guess of mine x3 It is a little dumb :P

OpenStudy (anonymous):

getting 2/3 on my quiz wouldn't be so horrible either x.x Although not as good as 3/3 hahah >.<

OpenStudy (anonymous):

There goes 1 and 2. Give me some times for the last one

OpenStudy (anonymous):

Thanks so much (: That looked like a whole lotta work!

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

I'm confused on the 2nd one o.o

OpenStudy (anonymous):

because me thinks you got -4,1,5 But that isn't one of the options Dx

OpenStudy (anonymous):

And I was like Gasp O.o Oh no !

OpenStudy (anonymous):

oh yea haha. sorry. I copied your stuff wrong. you want me to do it over?

OpenStudy (anonymous):

Aw :/ I feel really bad wasting your time like this )x

OpenStudy (anonymous):

is one option (-8,-7,5)

OpenStudy (anonymous):

yus (:

OpenStudy (anonymous):

I listed the options above in a comment haha for the third question too!

OpenStudy (anonymous):

I might have gotten different answers for the 3rd one

OpenStudy (anonymous):

sorry its having a tough time opening one sec haha

OpenStudy (anonymous):

just breezing through our work herer haha

OpenStudy (anonymous):

your work * lol xP

OpenStudy (anonymous):

I really appreciate your time, your awesome honestly, I must say!

OpenStudy (anonymous):

np. Im glad I helped.

OpenStudy (anonymous):

3 lbs peanuts, 6 lbs almonds, 2 lbs raisins : 8 lbs peanuts, 1 lb almonds, 2 lbs raisins : 6 lbs peanuts, 3 lbs almonds, 2 lbs raisins : 8 lbs peanuts, 2 lbs almonds, 1 lbs raisins So the 3rd one lol? your estimate?

OpenStudy (anonymous):

6 lbs of peanuts, 3 lbs of almonds, 2 lbs of raisin. 6+3+2=11 6(1.5)+3(3)+2(1.5)=21 and 6=3*2 I misunderstood the english thats all haha. it was not y=2x, it's x=2y...

OpenStudy (anonymous):

Alright (: Thanks a bunch ! I hope I can help you out sometime in the future haha! you deserve it.

OpenStudy (anonymous):

we'll see haha. Night... I gotta rest for tomorrow

OpenStudy (anonymous):

Goodnight (: